# ARTICLE TYPE PAGURI: a user experience study of creative interaction with text-to-music models

### v1

@@ Abstract
In recent years, text-to-music models have been the biggest breakthrough in automatic music generation. While they are unquestionably a showcase of technological progress, it is not clear yet how they can be realistically integrated into the artistic practice of musicians and music practitioners. This paper aims to address this question via Prompt Audio Generation User Research Investigation (PAGURI), a user experience study where we leverage recent text-to-music developments to study how musicians and practitioners interact with these systems, evaluating their satisfaction levels. We developed an online tool through which users can generate music samples and/or apply recently proposed personalization techniques, based on fine-tuning, to make the text-to-music model generate sounds closer to their needs and preferences. Using questionnaires, we analyzed how participants interacted with the proposed tool, to understand the effectiveness of text-to-music models in enhancing users’ creativity. Results show that even if the audio samples generated and their quality may not always meet user expectations, the majority of the participants would incorporate the tool in their creative process. Furthermore, they provided insights into potential enhancements for the system and its integration into their music practice.
@@
Up to six keywords (optional).
@@ 1. Introduction
We developed an interface through which the user can indicate through a text prompt the audio sample they wish to generate. Additionally, the user can upload up to desired audio samples to personalize the TTM generative model. For the study, we use AudioLDM2 (Liu et al., 2023b, ) as a generative model, and a TTM personalization technique proposed in (Plitsis et al.,, 2023) to let the users fine-tune the model according to the music samples of their choice. We then conducted a user experience study, both online and in presence, where users can interact with PAGURI and we analyzed their experience using questionnaires and open questions. Specifically, we first quantify their background related to music and AI tools, and then we analyze their level of satisfaction after each use of the TTM model, until reaching the desired result. Finally, we let them answer a questionnaire analyzing the whole experience and gathering information related to the perceived usability of the TTM model in music practice. We consider the findings significant contributions to the research community, serving as valuable guidelines for structuring user experience tests involving generative models. Additionally, they represent an initial step in fostering practical discussions regarding user interactions with such models, besides being guidelines for the development and evaluation of TTM models. Supplementary material containing the answers to all questions and feedback is available on the accompanying website111https://paguri-ismir-2024.github.io/PAGURI/.
@@ 2.1 Text-to-music models
Language models (Vaswani et al.,, 2017) provide extremely efficient capabilities in generating content with long-term context. As such, they have increasingly become the backbone of recently proposed generative models. Many works have applied them to the task of raw audio music and audio synthesis, employing different types of architectures. The first proposed model was AudioLM (Borsos et al.,, 2023), a multi-stage transformer-based model operating on tokens extracted from audio via SoundStream (Zeghidour et al.,, 2021), and from text via w2v-BERT (Chung et al.,, 2021). This methodology was further developed in MusicLM (Agostinelli et al.,, 2023), where the joint music-text embedding model MuLan (Huang et al.,, 2022) was applied to overcome data scarcity. Similar transformer-based approaches were proposed for general audio synthesis using auto-regressive AudioGen (Kreuk et al.,, 2022), MusicGen (Copet et al.,, 2024), as well as non-auto-regressive models such as MAGNeT (Ziv et al.,, 2024). Several diffusion-based models were also proposed. The first one, DiffSound (Yang et al.,, 2023) considered audio generation by proposing a text-conditioned diffusion model operating on tokenized audio. Later on, several diffusion-based approaches were proposed also for text-to-music generation, such as Make-an-Audio (Huang et al.,, 2023), AudioLDM (Liu et al., 2023a, ), and its evolution AudioLDM2 (Liu et al., 2023b, ).
Question | 1 | 2 | 3 | 4 | 5 |
1) I spend a lot of my free time doing music-related activities | 0 % | 8.3% | 25% | 41.7% | 25% |
2) I enjoy writing about music, for example on blogs and forums | 20.8 % | 54.2% | 16.7% | 8.3% | 0% |
3) If somebody starts singing a song I don’t know, I can usually join in | 12.5 % | 12.5% | 33.3% | 33.3% | 8.3% |
4) I can sing or play music from memory | 0 % | 4.2% | 8.3% | 41.7% | 45.8% |
5) I often read or search the internet for things related to music | 0 % | 0% | 25% | 16.7% | 58.3% |
6) I am able to identify what is special about a given musical piece | 4.2 % | 4.2% | 12.5% | 41.7% | 37.5% |
7) I often read or search the internet for things related to AI tools | 16.7 % | 20.8% | 29.2% | 29.2% | 4.2% |
8) AI can be a tool to support human activities | 0 % | 0% | 25% | 25% | 50% |
2.2 Personalization
@@
Personalization of a model involves adjusting pre-trained generative models to specific examples that the model has not encountered before, as they were not included in the original dataset. In the context of language-based generative models, the task was first introduced for images in (Gal et al.,, 2023), where the authors proposed the Textual-Inversion technique. This technique allows injecting into the model information from data that it did not encounter during training. This is performed by using a few images whose text-embedding representations, or pseudoword (i.e. a token), are learned in the embedding space of the text encoder. Anyway, this technique is limited by the expressive capabilities of the pre-trained generative model. To overcome this limitation, the DreamBooth approach was proposed in (Ruiz et al.,, 2023). In this case, the desired examples are represented using rare token identifiers, and, consequently, the considered generative model (usually diffusion-based) is fine-tuned to learn the desired content.
@@ 2.3 Interaction Studies
In the audio/music domain, previous research has explored the perception of AI-generated music (Chu et al.,, 2022) and the opportunities and challenges of AI for music creation from a human perception (Newman et al.,, 2023). Several interfaces have been proposed to enable the interaction between users and generative models (Simon et al.,, 2008; Huang et al.,, 2019; Louie et al.,, 2020; Rau et al.,, 2022; Zhang et al.,, 2021; Zhou et al.,, 2020, 2021; Yakura and Goto,, 2023). Among these, the work closest to the one proposed in this paper is IteraTTA (Yakura and Goto,, 2023), an interface for TTM generation enabling iterative exploration of prompts and audio priors. Differently from PAGURI, IteraTTA did not contain the possibility of personalizing the model using audio samples of choice. Moreover, our goals and design strategy are different. In IterATTA the objective was mainly to design an interface for supporting novice users, analyzing user behavior ex-post. In our study, instead, the simple PAGURI interface was designed to study how music practitioners interacted with these models. For this reason, the whole user experience test was conducted live with a more structured interview procedure based on questionnaires.
@@ 3.1 PAGURI Interface
The Prompt Audio Generator User Research Investigation (PAGURI) proposed in this paper presents an interface through which users can generate audio samples using state-of-the-art TTM models. Through the interface, users can generate audio samples giving textual input prompts of their choice to the generative model. They can specify the number of audio tracks to generate and their duration, for each iteration. Additionally, through the same interface, users can upload up to preferred audio files to personalize AudioLDM2. Subsequently, they can obtain a personalized model using the DreamBooth personalization technique as in (Plitsis et al.,, 2023). In this case, the user can specify an instance word used as an identifier to be attached to the new audio files, and an object class which indicates a more general description of the audio, which may indicate the instrument, genre, or musical style of the concepts. Each user can select to either personalize the off-the-shelf AudioLDM2 model or a previously personalized (by the same user) version of it. The fine-tuning duration can be chosen between Fast, Medium and Slow. Each of them applies the DreamBooth personalization procedure for different durations, ranging from a minimum of minutes (Fast) to a maximum of minutes per iteration (Slow). While longer training times enable to obtain better results, we constrained the maximum considered time to retain the interactivity aspect of the study. PAGURI interface has been developed using Python programming language, with the use of Jupyter Notebook and ipywidgets library222https://github.com/jupyter-widgets/ipywidgets. A detailed presentation and images showing PAGURI interface can be found on the supplementary material1.
3.2 Procedure
The procedure was the same for both in-person and online experiments. It can be divided into three substeps:
@@ 3.2 Procedure
Preliminary analysis. The participants were introduced to the study and they were required to answer a brief questionnaire containing demographic questions (age, nationality, etc.). To frame the knowledge of participants concerning TTM models, we asked them to answer a second questionnaire regarding their musical knowledge and experiences with AI tools, using questions partially taken from the Goldsmith Musical Sophistication Index (Gold-MSI) (Müllensiefen et al.,, 2014).
@@
Text-To-Music interaction. Participants were asked to interact with the PAGURI interface, generating audio samples and personalizing the TTM model to suit their preferences. After each generation iteration, participants were asked to complete the Model Evaluation Survey1, expressing their satisfaction with the generated audio samples regarding their consistency with respect to the input prompt, audio quality, and alignment with their general expectations.
@@ 4. Results
Question | 1 | 2 | 3 | 4 | 5 |
1) The generated audio(s) is consistent with respect to the input prompt | 18.8 % | 18.8% | 20.7% | 28.9% | 12.8% |
2) The audio(s) generated is consistent with respect to my expectations | 9.9 % | 16.4% | 27.3% | 28.9% | 17.4% |
3) The quality of the audio(s) generated is consistent with my expectations | 14.8 % | 19.1% | 20.7% | 28.6% | 16.8% |
@@ 4.1 Demographics Analysis of Participants
A total of people participated to the study, with an average age of years (sd = 5.01). identified as He/Him, while identified as She/Her. To guarantee anonymity, each participant has been assigned an ID. 95% of the participants were of Italian nationality. In terms of current occupation, 62.5% of the participants were master’s degree students, (mainly from the Master of Science in Music and Acoustic Engineering from Politecnico di Milano), 4.2% were bachelor’s degree students, while the remaining 33.3% were workers. Most participants are currently active in various musical endeavors: of them are part of a band or an orchestra, are music producers/mastering engineers or DJs, 2 of them are dance teachers and one is a record label manager. 62.5% of the participants conducted the experiments in person, while 37.5% participated remotely via the Zoom platform.
@@ 4.2 Participants Musical knowledge and AI Tool Experience: Demographics and Analysis
The first part of the experiment focused on understanding the relationship of participants with AI tools and music in general. To gather this insight, participants were asked to complete a questionnaire. In Table 1, we present questions where responses were rated on a Likert scale from 1 to 5. For other responses, we provide details in the text. Most of the participants have a strong interest in music, with listening to music daily for more than one hour. engaged in regular, daily practice of an instrument for more than years. At the peak of their interest, more than of the participants practiced for one hour or more daily. Among the participants, of them can play at least one instrument (among which guitar, piano, flute, drums, bass guitar, trombone, ocarina) or are singers.
Regarding the relationship between users and AI tools, is familiar with AI tools and the was able to mention some of them such as ChatGPT (Wu et al.,, 2023), Copilot (Bird et al.,, 2022), Dall-E (Ramesh et al.,, 2021), and Gemini (Team et al.,, 2023). Some TTM tools were also reported, such as Riffusion (Forsgren and Martiros,, 2022), and MusicGen (Agostinelli et al.,, 2023). Focusing on TTM tools, the is aware of their existence, although most of them were unable to explicitly name examples of TTM AI tools. However, while the majority of the participants regularly employ artificial intelligence tools, only the of participants use TTM models.
@@ 4.3 Text-To-Music interaction
In total, 304 textual prompts (mean = 12.6 prompts per user, sd = 4.56) were given to the TTM model, for a total of 1148 audio files generated (mean = 47.8 audio files per user). Additionally, there were 39 personalizations made by participants (mean = 1.63 per user), meaning 39 customizations of one of the AudioLDM2 models (either the original AudioLDM2 or a previously personalized model) were done using the DreamBooth technique (Plitsis et al.,, 2023; Ruiz et al.,, 2023). Different audio files were used as input to personalize the model: electric guitar riffs, one-shot bass or synth lead samples, and 8-10 second tracks of musical beats or songs, among others.
@@
Question | 1 | 2 | 3 | 4 | 5 |
1) I enjoyed the interaction with the text-to-music generation system | 0 % | 0% | 0% | 37.5% | 62.5% |
2)The data generated by the model are consistent with respect to the desired audio file provided by the user* | 4.5 % | 0.0% | 31.8% | 45.5% | 18.2% |
3) The waiting time of the fine-tuning of the model is proportionate with the quality of the generated audio* | 0.0 % | 9.1% | 18.2% | 40.9% | 31.8% |
4) The audio generated by the personalized model is better than the audio generated by the based model* | 4.5 % | 18.2% | 27.3% | 13.6% | 36.4% |
5) There is consistency between input prompt and audio(s) generated | 4.2 % | 4.2% | 33.3% | 58.3% | 0% |
6) The use of text-to-music models can support musicians in musical creation endeavours | 4.2 % | 4.2% | 20.8% | 29.2% | 41.7% |
7) I would use this system again | 4.2 % | 12.5% | 12.5% | 33.3% | 37.5% |
@@ 4.3.1 Personalization of the TTM model
Most of users personalized the TTM models giving audio samples as input. They all started from the off-the-shelf pretrained AudioLDM2 model.
In general, the users really appeciated the possibility of personalizing the model according to their taste. We here report some comments from the participants. If the code of the participant is followed by , the sentence has been directly translated from italian (this is valid for the whole paper). P4: ”It further improves the quality of the modeling the task of generating samples from a specific genre. Even if this means that the quality in very different genres is decreased it’s not a big deal, as we can still fall back to the original model or even fine-tune it again for a different task.” P7: ”It is good because it can fit better your own music tastes”. P18: ”The result is more likeable to the user.”
@@
However, issues concerning the copyright related to the TTM models personalization were also raised. P17 thinks that ”the riff generated after customization is too similar to that of the input data, at risk of plagiarism.”
Users were particulary impressed by the personalization ability of the models with respect to rhythms and percussive sounds. P3*: ”In terms of tone generation of the customization, the drums are better reproduced.” P10*: ”With the basic model, the audio generated does not have a rhythm or cadence, while with the customized model there is more rhythm and the tempo is better marked.” P14* noticed that during the personalization ”the sound is very dissonant (bad even on a harmonic level), but the percussive sounds are well reproduced.” P21*: ”It is very good at personalizing the rhythm.”.

### v2

@@ Abstract
In recent years, text-to-music models have been the biggest breakthrough in automatic music generation. While they are unquestionably a showcase of technological progress, it is not clear yet how they can be realistically integrated into the artistic practice of musicians and music practitioners. This paper aims to address this question via Prompt Audio Generation User Research Investigation (PAGURI), a user experience study where we leverage recent text-to-music developments to study how musicians and practitioners interact with these systems, evaluating their satisfaction levels. We developed an online tool through which users can generate music samples and/or apply recently proposed personalization techniques, based on fine-tuning, to allow the text-to-music model to generate sounds closer to their needs and preferences. Using questionnaires, we analyzed how participants interacted with the proposed tool, to understand the effectiveness of text-to-music models in enhancing users’ creativity. Results show that even if the audio samples generated and their quality may not always meet user expectations, the majority of the participants would incorporate the tool in their creative process. Furthermore, they provided insights into potential enhancements for the system and its integration into their music practice.
@@
Text-to-Music, Generative Models, Human-AI Interaction, Human-computer co-creativity.
@@ 1. Introduction
[…]
@@ 2.1 Text-to-music models
Language models (Vaswani et al.,, 2017) provide extremely efficient capabilities in generating content with long-term context. As such, they have increasingly become the backbone of recently proposed generative models. Many works have applied them to the task of raw audio music and audio synthesis, employing different types of architectures. The first proposed model was AudioLM (Borsos et al.,, 2023), a multi-stage transformer-based model operating on tokens extracted from audio via SoundStream (Zeghidour et al.,, 2021), and from text via w2v-BERT (Chung et al.,, 2021). This methodology was further developed in MusicLM (Agostinelli et al.,, 2023), where the joint music-text embedding model MuLan (Huang et al.,, 2022) was applied to overcome data scarcity. Similar transformer-based approaches were proposed for general audio synthesis using auto-regressive AudioGen (Kreuk et al.,, 2022), MusicGen (Copet et al.,, 2024), as well as non-auto-regressive models such as MAGNeT (Ziv et al.,, 2024). Several diffusion-based models were also proposed. DiffSound (Yang et al.,, 2023), the first one, considered audio generation by proposing a text-conditioned diffusion model operating on tokenized audio. Later on, several diffusion-based approaches were proposed also for text-to-music generation, such as Make-an-Audio (Huang et al.,, 2023), AudioLDM (Liu et al., 2023a, ) and its evolution AudioLDM2 (Liu et al., 2023b, ), MUSTANGO (Melechovsky et al.,, 2024), and Stable Audio Open (Evans et al.,, 2024). More recently, also several commercial solutions were proposed, such as Suno AI (Suno AI,, 2024), which raised 125 million $ in funding, and Udio (Udio,, 2024). These models became so popular that major record companies filed lawsuits against them, in order to regulate the use of copyrighted material for music generation.
2.2 Personalization techniques
@@
The personalization of a model involves adjusting pre-trained generative models to specific examples that the model has not encountered before, as they were not included in the original dataset. In the context of language-based generative models, the task was first introduced for images in (Gal et al.,, 2023), where the authors proposed the Textual-Inversion technique. This technique allows injecting into the model information from data that it did not encounter during training. This is performed by using a few images whose text-embedding representations, or pseudoword (i.e. a token), are learned in the embedding space of the text encoder. This technique is limited by the expressive capabilities of the pre-trained generative model. To overcome this limitation, the DreamBooth approach was proposed in (Ruiz et al.,, 2023). In this case, the desired examples are represented using rare token identifiers, and, consequently, the considered generative model (usually diffusion-based) is fine-tuned to learn the desired content.
@@ 2.3 Interaction Studies
In the audio/music domain, previous research has explored the perception of AI-generated music (Chu et al.,, 2022) and the opportunities and challenges of AI for music creation from a human perception (Newman et al.,, 2023). Several interfaces have been proposed to enable the interaction between users and generative models (Simon et al.,, 2008; Huang et al.,, 2019; Louie et al.,, 2020; Rau et al.,, 2022; Zhang et al.,, 2021; Zhou et al.,, 2020, 2021; Yakura and Goto,, 2023). Among these, the closest work to the one proposed in this paper is IteraTTA (Yakura and Goto,, 2023), an interface for TTM generation enabling iterative exploration of prompts and audio priors. Differently from PAGURI, IteraTTA did not contain the possibility of personalizing the model using audio samples of choice. Moreover, our goals and design strategy are different. The objective of IterATTA was mainly to design an interface for supporting novice users, analyzing user behavior ex-post. In our study, instead, the simple PAGURI interface was designed to study how music practitioners interacted with these models. For this reason, the whole user experience test was conducted live with a more structured interview procedure based on questionnaires.
@@ 3.1 PAGURI Interface
The Prompt Audio Generator User Research Investigation (PAGURI) proposed in this paper presents an interface through which users can generate audio samples using state-of-the-art TTM models, depicted in Fig. 1. Through the interface, users can generate audio samples giving textual input prompts of their choice to the generative model. They can specify the number of audio tracks to generate and their duration, for each iteration. Additionally, through the same interface, users can upload up to preferred audio files to personalize AudioLDM2. Since the number of files used for personalization is directly linked with the timing of the procedure, we empirically selected a number that allowed a reasonable tradeoff between speed (needed to preserve interactivity) and quality of the results. Subsequently, they can obtain a personalized model using the DreamBooth personalization technique as implemented in (Plitsis et al.,, 2023). In this case, the user can specify an instance word used as an identifier to be attached to the new audio files, and an object class which indicates a more general description of the audio, which may indicate the instrument, genre, or musical style of the concepts. Each user can select to either personalize the off-the-shelf AudioLDM2 model or a previously personalized (by the same user) version of it. The fine-tuning duration can be chosen between Fast, Medium and Slow. Each of them applies the DreamBooth personalization procedure for different durations, ranging from a minimum of minutes (Fast) to a maximum of minutes per iteration (Slow). While longer training times enable to obtain better results, we constrained the maximum considered time to retain the interactivity aspect of the study. PAGURI interface has been developed using Python programming language, with the use of Jupyter Notebook and ipywidgets library222https://github.com/jupyter-widgets/ipywidgets. A detailed presentation and images showing PAGURI interface can be found on the supplementary material1.
3.2 Experiment procedure
[…]
The whole procedure is divided into three steps:
@@ 3.2 Experiment procedure
Preliminary analysis. The participants were introduced to the study and they were required to answer a brief questionnaire containing demographic questions (age, nationality, etc.). To frame the knowledge of participants concerning TTM models, we asked them to answer a second questionnaire regarding their musical knowledge and experiences with AI tools, using questions partially taken from the Goldsmiths Musical Sophistication Index (Gold-MSI) (Müllensiefen et al.,, 2014).
@@
Text-To-Music interaction. The core part of the experiment. Participants were asked to interact with the PAGURI interface. At each iteration, participants had the option to input a new prompt into the TTM model, choosing whether or not to personalize the model with the audio of their choice, and then generate a desired number of new audio samples. After each generation iteration, participants were asked to complete the Model Evaluation Survey1, expressing their satisfaction with the generated audio samples regarding their consistency concerning the input prompt, audio quality, and alignment with their general expectations.
@@ 4.1 Demographics Analysis of Participants
[…]
@@ 4.2 Participants Musical Knowledge and AI Tool Experience: Demographics and Analysis
The first part of the experiment focused on understanding the relationship of participants with AI tools and music in general. To gather this insight, participants were asked to complete a questionnaire. In Fig 2, we display the subset of questions rated on a 5-point Likert scale, using a diverging bar chart. The chart illustrates the number of respondents for each answer, with positive responses on the right and negative responses on the left, centered around the midpoint. Neutral answers are split between positive and negative categories. The length of each bar represents the number of responses for each score, allowing easy comparison of the distribution. Each color indicates the group of responses corresponding to a specific score. For other responses, we provide details in the text.
Most of the participants have a strong interest in music, with listening to music daily for more than one hour. engaged in regular, daily practice of an instrument for more than years. At the peak of their interest, more than of the participants practiced for one hour or more daily. Among the participants, of them can play at least one instrument (among which guitar, piano, flute, drums, bass guitar, trombone, ocarina) or are singers.
Regarding the relationship between users and AI tools, are familiar with AI tools and were able to mention some of them such as ChatGPT (Wu et al.,, 2023), Copilot (Bird et al.,, 2022), Dall-E (Ramesh et al.,, 2021), and Gemini (Team et al.,, 2023). Some TTM tools were also reported, such as Riffusion (Forsgren and Martiros,, 2022), and MusicGen (Agostinelli et al.,, 2023). Focusing on TTM tools, are aware of their existence, although most of them were unable to explicitly name examples of TTM AI tools. However, while the majority of the participants regularly employ artificial intelligence tools, only the of participants use TTM models.
Iteration | Prompt |
1 | a song about summer containing an electric guitar lead on top of an ukulele rhythm and a glockenspiel. The glockenspiel is off |
2 | smooth jazz being played from the other room; the main line is played with a saxophone, backed by a leslie organ |
3 | An abcdef kalimba melody in a 90 BPM midtempo music drop |
4 | An abcdef voice singing a gregorian chant |
5 | A kalimba with a harsh bitcrusher applied to it |
6 | A abcdef kalimba with a bitcrusher applied to it |
Iteration | Prompt |
1 | A sound of an electric guitar |
2 | A flute playing ”My heart will go on” by Celine Dion |
3 | A ballad in the style of sas pirate metal |
4 | A jazz music in the style of sas |
5 | A doom metal music in the style of sas pirate metal |
6 | A 90’s disco hit in the style of sas pirate metal |
7 | Vegeta singing a song in the style of sas pirate metal |
8 | ”Epic sax guy” in the style of sas pirate metal |
9 | A dog barking in the style of sas pirate metal |
10 | A dog barking in the style of pirate metal |
11 | A rubber chicken singing a song in the style of sas pirate metal |
12 | a sas pirate metal song, but with the lyrics made by a door bell ring |
@@ 4.3 Text-To-Music interaction
In total, 304 textual prompts (mean = 12.6 prompts per user, s.d. = 4.56) were given to the TTM model, for a total of 1148 audio files generated (mean = 47.8 audio files per user. Multiple audio files were generated per each prompt). Additionally, there were 38 personalizations made by participants (mean = 1.58 per user, s.d. = 0.86), meaning 38 customizations of one of the AudioLDM2 models (either the original AudioLDM2 or a previously personalized model) were done using the DreamBooth technique (Plitsis et al.,, 2023; Ruiz et al.,, 2023). Different audio files were used as input to personalize the model: electric guitar riffs, one-shot bass or synth lead samples, and 8-10 second tracks of musical beats or songs, among others.
@@ 4.3.1 Personalization of the TTM model
Most users personalized the TTM models giving audio samples as input. They all started from the off-the-shelf pretrained AudioLDM2 model.
In general, the users really appreciated the possibility of personalizing the model according to their taste. We here report some comments from the participants. If the code of the participant is followed by , the sentence has been directly translated from Italian (this is valid for the whole paper). P4: ”It further improves the quality of the modeling the task of generating samples from a specific genre. Even if this means that the quality in very different genres is decreased it’s not a big deal, as we can still fall back to the original model or even fine-tune it again for a different task.” P7: ”It is good because it can fit better your own music tastes”. P18: ”The result is more likeable to the user.”
@@
However, issues concerning the copyright related to the TTM model’s personalization were also raised. P17 thinks that ”the riff generated after customization is too similar to that of the input data, at risk of plagiarism.”
Users were particularly impressed by the personalization ability of the models with respect to rhythms and percussive sounds. P3*: ”In terms of tone generation of the customization, the drums are better reproduced.” P10*: ”With the basic model, the audio generated does not have a rhythm or cadence, while with the customized model there is more rhythm and the tempo is better marked.” P14* noticed that during the personalization ”the sound is very dissonant (bad even on a harmonic level), but the percussive sounds are well reproduced.” P21*: ”It is very good at personalizing the rhythm.”.